Re: Shift report 20011014 00:00-08:00

From: Ramiro Debbe (debbe@sgs1.hirg.bnl.goV)
Date: Sun Oct 14 2001 - 10:40:52 EDT

  • Next message: Apache: "Shift report 20011014 08:00-16:00"

    T2 A has a leakeage current of some 800 or 900 nA at 1260 with no beam in
    the IR. The current should go up when there are collisions. Currents
    around 1 micro amp are to be expected. I strongly
    disagree to any change in anode HV in the TPCs, we know the TPCs are
    working well, lets not start guessing solutions to imaginary problems.
    Next time somebody feels an urge to change some TPC parameter, please let
    me know first.
    
    Ramiro 
    
    On Sun, 14 Oct 2001, Apache wrote:
    
    > 
    > ------------------------------------------------------------
    > Date: Sun Oct 14 08:13:48 2001, Shift: 00:00-08:00
    > Supervisor: J.I.Jordre, with: 
    > Log book#9 pages 149 - 151
    > 
    > Shift Summary:
    > ------------------------------------------------------------
    > 1. Beam from 01:24 to 05:14 and again from 06:48 which is still going strong.
    >    ZDC rates comparable to the other experiments but barely over 200Hz at most.
    > 
    > 2. (MRS: 4B @ 45 deg., FS: 1/5A @ 12 deg.)
    >    Run 5318	297k trigger 1 events
    >    Run 5319	130k trigger 1 events
    >    Run 5324	180k trigger 1 events by 08:00
    >    DVM, TPC pedestal, DVM and pedestal runs 5317, 5320, 5321 and 5322.
    > 
    > 3. T2 anodes A had ~1.05uA current at beginning of stores with default 1260V. 
    >    I therefore chose to lower the anode A voltage by 20V to get down to the safer
    >    0.85uA current. Looking at SuperMonitor, TED and TPC3D I could not gauge any
    >    loss in efficiency. The distributions of max ADC in the front part of T2 is
    >    more or less the same as the rear part of it. Therefore runs 5318 and 5324 are 
    >    taken with T2 anodes A at 1240V.
    >    ZDC blue high gain ADC 1's pedestal is no good, basically just showing up as 
    >    1s in SuperMonitor plot.
    >    H1 calib. slat 1's ADC distribution is bad. 
    >    Problem with TOFW slat 92 reported by YKL is in TDC only, not ADC.
    >    BBR 25 has low gain.
    >    Tile 34 still not working.
    >    TOFW calib. slat 3 top still has wide pedestals, as reported by YKL.
    >    
    > ------------------------------------------------------------
    > 
    



    This archive was generated by hypermail 2b30 : Sun Oct 14 2001 - 10:40:04 EDT